# X2I: Seamless Integration of Multimodal Understanding into Diffusion Transformer via Attention Distillation

### v1

@@ 1 Introduction
The text encoder in T2I is critical for generating semantically relevant images, as seen from early models such as CLIP [58] to later models such as T5 [59] and advanced LLMs [73, 13]. The main goal is to improve the semantic understanding of text without considering the impact of other modalities on visual output. In contrast, VLMs, including early approaches like Flamingo [2] and BLIP2 [35], along with newer models like Llama-Vision [12], QwenVL [3], and InternVL [10], have effectively integrated pre-trained visual encoders with LLMs for enhanced text and visual comprehension. Is there a more streamlined way to transfer VLM or MLLM capabilities to DiT to enable diverse modal inputs? The industry is primarily tackling this challenge through two strategies. One method, similar to PEA-Diffusion [46], utilizes feature distillation by aligning multilingual LLM encoders to the original dimensions using MLP. However, it requires substantial resources of 1600 GPU hours and large-scale image-text datasets. This distillation relies on the U-Net[64] block’s outputs and limits multimodal input support. Another method is the VLMs-based replacement of T2I [44], but it is costly, requiring extensive training or fine-tune of T2I or MLLMs, necessitating substantial data quality and quantity [69, 22]. Currently, image generation models lack multimodal understanding beyond text and vision.

### v2

@@ 1 Introduction
The text encoder in T2I is critical for generating semantically relevant images, as seen from early models such as CLIP [59] to later models such as T5 [60] and advanced LLMs [74, 13]. The main goal is to improve the semantic understanding of text without considering the impact of other modalities on visual output. In contrast, VLMs, including early approaches like Flamingo [2] and BLIP2 [35], along with newer models like Llama-Vision [12], QwenVL [3], and InternVL [10], have effectively integrated pre-trained visual encoders with LLMs for enhanced text and visual comprehension. Is there a more streamlined way to transfer VLM or MLLM capabilities to DiT to enable diverse modal inputs? The industry is primarily tackling this challenge through two strategies. One method, similar to PEA-Diffusion [46], utilizes feature distillation by aligning multilingual LLM encoders to the original dimensions using MLP. However, it requires substantial resources of 1600 GPU hours and large-scale image-text datasets. This distillation relies on the U-Net[65] block’s outputs and limits multimodal input support. Another method is the VLMs-based replacement of T2I [44], but it is costly, requiring extensive training or fine-tune of T2I or MLLMs, necessitating substantial data quality and quantity [70, 22]. Alternatively, a solution akin to GlueGen[57] could be employed to align the features of a single-modal or multimodal encoder with the latent space of existing T2I models. However, in terms of strategy, aligning solely through the encoder still results in suboptimal overall performance.
@@ 4.5 X2I Training Convergence
4.6 X2I’s reasoning ability and multi-turn dialogue generation ability
MLLMs themselves possess certain reasoning abilities and multi-turn dialogue capabilities. X2I also inherits some of these related abilities. Fig. 5 illustrates the reasoning abilities based on text as well as image-based reasoning abilities. It is important to note here that the features fed into X2I originate from the characteristics of MLLMs’ answers. Fig. 6 demonstrates X2I’s multi-turn dialogue capabilities, indicating that the generated images exhibit a certain degree of coherence and fidelity. Although these capabilities only represent a small portion of the abilities inherited from MLLMs, they provide more possibilities for improving the productivity of images.
Figure 5: X2I’s reasoning ability.
Figure 6: X2I’s multi-turn dialogue generation ability.